OpenStudy (anonymous):

If 6.0 moles of X and 5.0 moles of Y react according to the reaction below, how many moles of the excess reactant will be left over at the end of the reaction? 3 X + 2Y -> X3Y2 1.0 mol Y left over 1.5 mol Y left over 4.0 mol X left over 2.5 mol X left over

OpenStudy (anonymous):

X + 2 Y = XY2 moles X = 3.0 the ratio between X and Y is 1 : 2 moles Y required = 3.0 x 2 = 6.0 because moles Y = 5.0 this means that Y is the limiting reactant moles X neede to react with 5.0 moles of Y == 5.0/2 = 2.5 moles X in excess = 3.0 - 2.5 = 0.5 0.5 mol X left over